Suppose that the market demand and supply equations in a perfectly competitive market are QD = 16 − 4P and QS = −2 + 2P, respectively. What is the full economic price if the government imposes a price ceiling of $2?
Answer #1

Market equilibrium when Qd=Qs

16-4P=-2+2P

18=6P

P=3 and Q=-2+6=4

when govt impose a price ceiling below equilibrium price then it is binding.

Thus full economic price=2